OpenStudy (anonymous):

4(2x − 3) = 20 solve please show steps.

jhonyy9 (jhonyy9):

multiplie term by term

jhonyy9 (jhonyy9):

4*2x = 4*(-3) =

OpenStudy (anonymous):

i thought it would be like this 8x - 12

Parth (parthkohli):

Distribute. \(\Large \color{MidnightBlue}{\Rightarrow 8x - 12 = 20 }\) Add 12 to both sides. \(\Large \color{MidnightBlue}{\Rightarrow 8x = 32 }\) Divide both sides by 8 \(\Large \color{MidnightBlue}{\Rightarrow x = 4 }\)

OpenStudy (anonymous):

Thanks everyone I'm bad with math Haha!

OpenStudy (anonymous):

ya but your an awesome DJ lol ;)